) A father said to his son, “I was as old as you are at present at time of your birth”. If the father’s age is 46 years now, the son’s age five years back was?
Answer by mamiya(56)

A father said to his son, “I was as old as you are at present at time of your birth”. If the father’s age is 46 years now, the son’s age five years back was?
let f be the father's age now and s the son's age now.
I was as old as you are at present at time of your birth, means s years ago the father was s. so f-s=s
f=2s
so now, the son is s=f/2 years old.
the son's age five years back was s-5= f/2 - 5
= 46/2 - 5
= 18
so the answer is 18